# MCA-Bench: A Multimodal Benchmark for Evaluating CAPTCHA Robustness Against VLM-based Attacks

### v2

@@ Abstract
As automated attack techniques rapidly advance, CAPTCHAs remain a critical defense mechanism against malicious bots. However, existing CAPTCHA schemes encompass a diverse range of modalities—from static distorted text and obfuscated images to interactive clicks, sliding puzzles, and logic-based questions—yet the community still lacks a unified, large-scale, multimodal benchmark to rigorously evaluate their security robustness. To address this gap, we introduce MCA-Bench, a comprehensive and reproducible benchmarking suite that integrates heterogeneous CAPTCHA types into a single evaluation protocol. Leveraging a shared vision–language model backbone, we fine-tune specialized cracking agents for each CAPTCHA category, enabling consistent, cross-modal assessments. Extensive experiments reveal that MCA-Bench effectively maps the vulnerability spectrum of modern CAPTCHA designs under varied attack settings, and—crucially—offers the first quantitative analysis of how challenge complexity, interaction depth, and model solvability interrelate. Based on these findings, we propose three actionable design principles and identify key open challenges, laying the groundwork for systematic CAPTCHA hardening, fair benchmarking, and broader community collaboration. Datasets are available at https://www.kaggle.com/datasets/luffy798/mca-benchmultimodal-captchas. Code is released at https://github.com/noheadwuzonglin/MCA-Bench.
@@
Figure 1: Data samples from MCA-Bench includes four categories and 20 sub-clusters of Point-and-Click Localization, Static Visual Recognition, Textual Logic Q&A and Interactive Manipulation.
1 Introduction
Type Dataset Image size #Images #Train Original image Single image
Static visual recognition Classic character CAPTCHA 160×60 5000 4800
Distorted-word CAPTCHA 800×200 5000 4800
Arithmetic-character CAPTCHA 130×48 5000 4800
Point-and-click localization 3×3 grid selection 672×672 5500 5300
Inverted-letter selection 224×224 5000 4800
Geometric-shape recognition 224×224 5898 5698
Brightness discrimination 224×224 5377 5177
Hollow-pattern recognition 224×224 5000 4800
Sequential-letter ordering 300×150 4999 4799
Full-image grid selection 512×512 5015 4815
Color discrimination 224×224 5000 4800
Vowel selection 224×224 5322 5122
Arithmetic-selection 240×240 5065 4865
Rotated-letter selection 224×224 5015 4815
3×3 jigsaw-swap selection 224×224 6505 6305
Interactive manipulation Sliding-block 600×360 5012 4812
Rotation-block 600×360 5041 4841
Alignment sliders 600×360 5006 4806
Textual logic Q&A Text-based arithmetic 1000 800
Commonsense reasoning 1000 800
@@ 1 Introduction
CAPTCHA (Completely Automated Public Turing test to tell Computers and Humans Apart) has long played a vital role in protecting online services from automated attacks. In recent years, growing research interest in AI security topics such as adversarial attacks van2023anti ; song2024anti ; song2025idprotector and digital watermarking ci2024ringid ; ci2024wmadapter ; liu2024image ; yang2024can has also drawn attention to the vulnerability of CAPTCHA systems. However, with rapid advances in deep learning, computer vision, and multimodal pretraining models, many CAPTCHA types once considered secure have now become susceptible to machine-learning-based attacksbursztein2011text ; sivakorn2016robot ; ye2018yet . Techniques such as GANs, vision-language models (VLMs), and reinforcement learning have enabled attackers to mimic human behavior with increasing precisiongoodfellow2014generative ; radford2021learning ; li2022blip ; shah2023lm . As a result, researchers have begun developing multimodal CAPTCHA datasets and evaluation frameworks to assess model performance across various CAPTCHA typesacien2020becaptcha ; acien2021becaptcha ; gao2021research . This makes it vital to reassess CAPTCHA’s real-world security to ensure a trustworthy Internet service ecosystemnoury2020deep ; gao2021research .
Despite the emergence of new interactive CAPTCHA forms, existing studies often focus on specific categories (e.g., text recognition, image selection) without offering comprehensive comparisonsbursztein2011text ; ding2025illusioncaptcha ; hernandez2021basecass . Moreover, there is no large-scale, multimodal benchmark for evaluating CAPTCHA designs against intelligent-agent attackshossen2022aaecaptcha ; searles2023empirical ; deng2024oedipus . This gap limits the identification of vulnerabilities and the development of more robust CAPTCHA systems. Therefore, a unified, standardized evaluation platform is urgently needed for security analysis across modalities and tasks.
@@
MCA-Bench is the first end-to-end CAPTCHA security benchmark spanning four modalities—static visual recognition, point-and-click localization, interactive manipulation and textual logic Q&A—across twenty real-world tasks. It provides over 180000 training samples and a 4000-item test set (200 examples per task), organized into four clusters that respectively evaluate OCR robustness to visual noise, target retrieval in complex scenes, human-like interaction behaviors, and multi-step language reasoning. Representative samples from the MCA-Bench dataset are shown in Figure 1.
@@
We use QwenVL-2.5-7B as the vision-language backbone, fine-tuned with LoRA adapters for each task. Training for static and logic CAPTCHAs is supervised with target labels, while for interactive tasks, human demonstration data is used for behavior cloning. A specially designed JSON protocol facilitates large-scale evaluation and integration.
@@
The evaluation framework uses pass rates as the core metric. Experiments show that multimodal VLMs achieve over 96% accuracy on simple tasks, while more complex ones (requiring physical interaction or multi-step logic) still pose significant challenges, with success rates as low as 2.5%. This highlights the coupling of ’visual confusion, depth of interaction, and semantic complexity’ with attack difficulty, providing actionable insights for future CAPTCHA design. MCA-Bench is released open-source, enabling easy replication and fostering ongoing cycles of attack, defense, and benchmark iteration. The main contributions are as follows.
@@
Table 2: Comparison of pass rates for various CAPTCHA types. This table presents the pass rates for different CAPTCHA task types, comparing the performance of the QwenVL-2.5-7B-Instruct model with human pass rates.
CAPTCHA Type QwenVL-2.5-7B-Instruct Pass Rate Human Pass Rate
Point-and-Click Localization 3×3 grid selection 0.96 0.88
Inverted-letter selection 0.52 0.94
Geometric-shape recognition 0.96 0.98
Brightness discrimination 0.665 0.78
Hollow-pattern recognition 0.995 0.98
Sequential-letter ordering 0.925 0.98
Full-image grid selection 0.35 0.74
Color discrimination 0.99 0.88
Vowel selection 0.975 0.80
Arithmetic selection 0.025 0.78
Rotated-letter selection 0.335 0.74
3×3 jigsaw-swap selection 0.96 0.80
Static Visual Recognition Classic character CAPTCHA 0.32 0.92
Distorted-word CAPTCHA 0.985 0.84
Arithmetic-character CAPTCHA 0.695 0.98
Textual Logic Q&A Text-based arithmetic 0.985 0.97
Commonsense reasoning 0.97 0.86
Interactive Manipulation Sliding-block 0.36 0.74
Rotation-block 0.28 0.76
Alignment sliders 0.55 0.72
2 Related Work
2.1 VLMs in Structured Image Tasks
@@
Visual Language Models (VLMs) bridge images and text through unified embeddings, with early methods like CLIP radford2021learning and ALIGN pmlr-v139-jia21b excelling in simple visual tasks but falling short in complex reasoning. Recent models integrate large language models (LLMs) with visual encoders to enhance cross-modal understanding—BLIP-2 li2022blip uses frozen components for efficient training, while MiniGPT-4 DBLP:conf/iclr/Zhu0SLE24 maps visual features to LLMs for stronger image-text QA. As research shifts toward structured tasks like tables and documents, which resemble CAPTCHA images with dense text and noise, models such as Qwen-VL wang2024qwen2vlenhancingvisionlanguagemodels improve encoder design and pretraining for better layout and multilingual comprehension. Architectures like LLaVA NEURIPS2023_6dcf277e and MiniGPT-4 DBLP:conf/iclr/Zhu0SLE24 adopt a unified visual projection and instruction-tuning framework, achieving robustness on distorted and occluded inputs. However, most VLMs lack CAPTCHA-specific adaptation, limiting their accuracy and generalization.
2.2 Evolution of Intelligent Agents Toward AGI
@@
An agent in AI refers to an autonomous entity that perceives its environment via sensors, makes decisions, and acts through actuators russell2021modern , characterized by autonomy, reactivity, and social interactivity—key traits driving progress toward Artificial General Intelligence (AGI) 10.1007/3-540-58095-6_9 ; 109c6f85f2734b649d5f7b7a1946b2c1 . Agent research has evolved from early symbolic agents with interpretable rule-based reasoning 10.1145/360018.360022 ; 10.5555/2974989 , yet limited in handling uncertainty and scalability 10.5555/110787 , to reactive agents offering fast perception-action loops 10.5555/1625995.1626091 , albeit with limited decision-making. The rise of reinforcement learning (RL) enabled agents like AlphaGo 44806 to learn from interactions, though challenges like sample inefficiency and poor generalization remained 10.1023/A:1015008417172 ; 1996cs……..5103K ; 1571980075547310080 ; 1573668924277769344 . Transfer learning and meta-learning improved task adaptation via knowledge reuse and learning-to-learn paradigms 10.5555/2772879.2772905 ; 2015arXiv151106342P ; 10.1109/TPAMI.2023.3292075 ; 2016arXiv161102779D ; 10.5555/3305381.3305498 ; 10.5555/3327345.3327436 ; rakelly*2019efficient ; Fakoor2020 , despite issues in task diversity and pre-training costs farebrother2019generalization . Recently, LLM-based agents have transformed the field, exhibiting emergent reasoning and planning abilitiesmckinney2020training ; achiam2023gpt ; wei2022emergent ; mann2020language , multimodal understanding (e.g., BLIP-2 li2023blip ), dynamic task decomposition (e.g., Voyager wang2023voyager ), and tool usage (e.g., Toolformer schick2023toolformer ), enabling general-purpose intelligence park2023generative ; liu2023training ; sumers2024cognitivearchitectureslanguageagents ; Wang_2024 . Their capabilities in zero-shot generalization wei2021finetuned ; sanh2021multitask ; chung2022scalinginstructionfinetunedlanguagemodels and social collaborationhong2024metagptmetaprogrammingmultiagent mark a paradigm shift in agent research.
2.3 Advances and Challenges in CAPTCHA Security
Early text CAPTCHAs used heavy distortion and noise, but segmentation flaws and CNN-based attacks soon achieved high recognition rates—even breaking reCAPTCHA under laboratory conditions chellapilla2005building ; simard2003best ; bursztein2011text . Image-classification schemes such as ASIRRA shifted effort to visual cognition, yet SVMs trained on public datasets ultimately compromised them elson2007asirra ; golle2008machine . Contemporary defences leverage deep object-detection networks alongside generative techniques (e.g., diffusion models and style-transfer) to create more diverse and intricate puzzles nian2022deep ; xing2023review ; jiang2023diff ; ding2025illusioncaptcha ; goodfellow2014generative ; luo2024image . Interactive CAPTCHAs, exemplified by reCAPTCHA v2, further blend visual tasks with user gestures such as drag-and-drop or click behaviour to raise the bar for automation shet2014you . Despite these advances, CAPTCHA schemes still struggle with adaptive difficulty tuning, behaviour consistency across devices, adversarial robustness, and—critically—the absence of standardized, reproducible benchmarks for end-to-end evaluationkumar2021benchmarks ; alsuhibany2016benchmark .
@@ 2.3 Advances and Challenges in CAPTCHA Security
Figure 2: Schematic overview of the MCA-Bench data-acquisition and annotation workflow. The four grey panels—from left to right—represent the dataset categories Static Visual Recognition, Interactive Manipulation, Point-and-Click Localization, and Textual Logic Q&A. The red labels (SVT, MIM-RU, CPTL, and LLM-GFP) denote the dedicated data-collection pipelines associated with each category. Each pipeline follows a unified four-stage paradigm: (i) determining the raw input format; (ii) applying task-specific geometric transformations, coordinate projections, or prompt-template generation; (iii) separating fine-grained annotation types; and (iv) saving the annotations to text files.
3 Dataset Construction Process
@@
This section outlines the MCA Bench pipeline—from raw sample collection to final benchmark release—while detailing the data-processing and annotation strategies for its four task clusters and twenty subtasks. The complete workflow appears in Figure 2.
3.1 Data Collection Sources and Compliance Strategies
@@
In developing the MCA-Bench benchmark suite, we implemented a comprehensive data collection and compliance management system to uphold scientific integrity and ensure legal and ethical standards in evaluating CAPTCHA security. Our approach combines independently created data and reused public datasets. For the former, we designed a diverse CAPTCHA dataset, including distorted text, obfuscated graphics, interactive tasks, and puzzles, based on a thorough analysis of technology trends and security needs. These designs are original, validated, and proprietary to our team.
For reused datasets, we carefully selected authorized academic, open-source, and industry resources, adhering to licensing agreements and privacy regulations. We ensured compliance through preprocessing steps like anonymization and mitigating privacy risks. Additionally, we established a multi-level review system with legal and ethics oversight, encrypted storage, and tiered access control to guarantee ongoing compliance and data security for MCA-Bench.
3.2 Data Collection and Processing
3.2.1 Text-based Task Data Collection
@@
For text-based CAPTCHA tasks, we developed a semi-automated LLM-driven pipeline to efficiently generate and filter math and commonsense questions, reducing manual effort while maintaining quality and diversity. By using adaptive prompts, we guided the Qwen LLM to create structured, semantically relevant question-answer pairs, incorporating knowledge constraints, task-guided sampling, and type-control parameters for difficulty and logic. The output was further refined through manual filtering with a custom evaluation protocol, assessing grammar, reasoning, and ambiguity to ensure robustness and clarity for real-world use.
3.2.2 Click-based Coordinate Task Data Collection
@@
We propose CPTL (Click-based Positioning and Target Localization), a multimodal benchmark designed to evaluate models’ spatial localization accuracy and image-language instruction alignment across varying scenario complexities. The framework consists of two main tasks: Free-form Pointing and Grid-based Discrete Pointing. In the Free-form Pointing task, we use a highly controllable image generation process combining procedural background perturbations with public visual datasets, creating images layered with semantic content, based on Flickr scenic photos. The Grid-based Discrete Pointing task employs a standardized 3×3 grid layout to assess model decision-making within constrained regions, dividing each image into 9 segments (0 to 8) to form a discrete click space. The image content is derived from a 64-class animal classification datasetanthony_therrien_2024 , Flickr scenic backgrounds, Google ReCaptcha V2 challenges, and manually designed targets.
@@
CPTL supports systematic, adjustable evaluation of spatial complexity and instruction modalities, enabling isolated assessment of spatial parsing and language comprehension. Its key advantage lies in simulating real-world CAPTCHA click scenarios while maintaining robust model performance under noisy backgrounds, multi-target interference, and weak semantic prompts. Experimental results demonstrate that multimodal models fine-tuned with CPTL exhibit strong localization and semantic understanding, providing a reliable evaluation baseline and optimization direction for future CAPTCHA security research.
3.2.3 Static Visual Recognition Task Data Collection
We propose SVT (Static Visual Textual Understanding), a framework designed to evaluate multimodal models’ ability to understand and reconstruct text from distorted images. SVT combines procedural image generation with language modeling constraints, allowing high controllability and complexity. It uses character-level geometric perturbations (e.g., affine transformations, warping, and perspective distortion) to generate challenging yet recognizable character patterns, while a stochastic interference generator adds random noise and occlusions. SVT effectively tests models’ character-level attention and resilience to structural noise, providing a comprehensive evaluation from visual perception to symbolic reasoning. Results show that SVT reveals bottlenecks in fine-grained character perception and serves as a valuable benchmark for assessing the robustness of pretrained multimodal models.
3.2.4 Interactive Behavior Task Data Collection
@@
We present a multimodal interaction modeling framework for interactive CAPTCHA tasks, based on real user operations, to enhance vision-language models’ dynamic understanding capabilities. Focusing on continuous motion interactions—such as sliding alignment, rotation calibration, and trajectory restoration—the framework requires models to perceive spatial structures, motion directions, and behavioral patterns. Data is collected through large-scale manual sampling across devices, capturing thousands of diverse user trajectories with timestamps and velocity, enabling realistic spatiotemporal behavior modeling.
@@
Tasks are designed around target state restoration, formalized into standardized sub-tasks with defined start, goal, and intermediate states. Interaction trajectories are serialized into behavioral vector streams and jointly encoded with image content to guide visual-behavior learning. Fine-tuning with high-quality samples improves the model’s decision consistency and action alignment. To evaluate performance, we introduce metrics including Center Deviation Error, Angular Restoration Accuracy, Slide Path Alignment Rate, and Motion Variability Index, offering comprehensive assessment beyond traditional image recognition benchmarks.
3.3 Data Annotation Strategy

### v6

@@ Abstract
As automated attack techniques rapidly advance, CAPTCHAs remain a critical defense mechanism against malicious bots. However, existing CAPTCHA schemes encompass a diverse range of modalities—from static distorted text and obfuscated images to interactive clicks, sliding puzzles, and logic-based questions—yet the community still lacks a unified, large-scale, multimodal benchmark to rigorously evaluate their security robustness. To address this gap, we introduce MCA-Bench, a comprehensive and reproducible benchmarking suite that integrates heterogeneous CAPTCHA types into a single evaluation protocol. Leveraging a shared vision–language model backbone, we fine-tune specialized cracking agents for each CAPTCHA category, enabling consistent, cross-modal assessments. Extensive experiments reveal that MCA-Bench effectively maps the vulnerability spectrum of modern CAPTCHA designs under varied attack settings, and—crucially—offers the first quantitative analysis of how challenge complexity, interaction depth, and model solvability interrelate. Based on these findings, we propose three actionable design principles and identify key open challenges, laying the groundwork for systematic CAPTCHA hardening, fair benchmarking, and broader community collaboration.
Codehttps://github.com/noheadwuzonglin/MCA-Bench
Datasetshttps://www.kaggle.com/datasets/luffy798/mca-benchmultimodal-captchas
@@
Figure 1: Data samples from MCA-Bench. Includes four categories and 20 sub-clusters of Point-and-Click Localization, Static Visual Recognition, Textual Logic Q&A and Interactive Manipulation.
Introduction
@@
CAPTCHA (Completely Automated Public Turing test to tell Computers and Humans Apart) has been crucial in protecting online services from automated attacks. However, with advancements in deep learning, computer vision, and multimodal pretraining models, many CAPTCHA types once considered secure are now vulnerable to machine-learning-based attacks (bursztein2011text; 10.1007/3-540-58095-6_9; wang2023voyager). Techniques such as GANs, vision-language models (VLMs), and reinforcement learning have enabled attackers to mimic human behavior with increasing precision (10.5555/2974989; noury2020deep; kumar2021benchmarks; schick2023toolformer). As a result, researchers have begun developing multimodal CAPTCHA datasets and evaluation frameworks to assess model performance across various CAPTCHA types (acien2020becaptcha; farebrother2019generalization; acien2021becaptcha). This makes it vital to reassess CAPTCHA’s real-world security to ensure a trustworthy Internet service (10.1145/360018.360022; farebrother2019generalization).
Existing studies often target specific CAPTCHA types without broad comparisons (bursztein2011text; ci2024ringid; 10.5555/3327345.3327436). The absence of a large-scale, multimodal benchmark (hernandez2021basecass; sanh2021multitask; ci2024wmadapter) limits systematic evaluation and hinders robust CAPTCHA design. A unified evaluation platform is urgently needed.
@@
MCA-Bench is the first end-to-end CAPTCHA security benchmark spanning four modalities—static visual recognition, point-and-click localization, interactive manipulation and textual logic Q&A—across twenty real-world tasks. It provides over 180000 training samples and a 4000-item test set, organized into four clusters that respectively evaluate OCR robustness to visual noise, target retrieval in complex scenes, human-like interaction behaviors, and multi-step language reasoning. Representative samples from the MCA-Bench dataset are shown in Figure 1.
@@
We use Qwen2.5-VL-7B as the vision-language backbone, fine-tuned with LoRA adapters for each task (hu2022lora). Training for static and logic CAPTCHAs is supervised with target labels, while for interactive tasks, human demonstration data is used for behavior cloning. A specially designed JSON protocol facilitates large-scale evaluation and integration.
@@
We use pass rate as the core metric. Evaluation shows multimodal VLMs exceed 96% accuracy on simple tasks but fall to as low as 2.5% on complex ones requiring physical interaction or multi-step reasoning. This reveals that visual confusion, interaction depth, and semantic complexity jointly drive attack difficulty, offering practical guidance for CAPTCHA design. MCA-Bench is open-sourced to enable reproduction and sustain iterative attack/defense benchmarking. The main contributions are as follows:
@@
Related Work
VLMs in Structured Image Tasks
@@
Visual Language Models (VLMs) unify image and text representations, with early models like CLIP (noury2020deep) and ALIGN (hossen2022aaecaptcha) performing well on simple tasks but lacking complex reasoning. Recent approaches integrate LLMs with visual encoders to enhance cross-modal understanding—e.g., BLIP-2 (kumar2021benchmarks) employs frozen components for efficiency, while MiniGPT-4 (Wang_2024) maps visual features to LLMs for improved QA. As focus shifts to structured inputs like tables and documents—similar to text-heavy, noisy CAPTCHA images—models such as Qwen-VL (song2025idprotector) advance encoder design and multilingual layout understanding. Architectures like LLaVA (li2023blip) and MiniGPT-4 leverage unified visual projection and instruction tuning to handle distortions and occlusions. However, most VLMs remain unadapted to CAPTCHA tasks, limiting their performance and generalization.
Evolution of Intelligent Agents Toward AGI
@@
An AI agent is an autonomous entity that perceives, decides, and acts in its environment (mckinney2020training), characterized by autonomy, reactivity, and social interactivity—traits crucial to AGI development (searles2023empirical; hong2024metagptmetaprogrammingmultiagent). Agent research has evolved from symbolic rule-based systems (luo2024image; 10.5555/3305381.3305498), limited by uncertainty and scalability (radford2021learning), to reactive agents with fast perception-action loops (10.5555/110787) but limited planning. Reinforcement learning enabled agents like AlphaGo (shah2023lm), though sample inefficiency and poor generalization persisted (park2023generative; pmlr-v139-jia21b). Transfer and meta-learning improved adaptation via knowledge reuse (10.5555/2772879.2772905; nian2022deep; deng2024oedipus; Fakoor2020), despite high pre-training costs (elson2007asirra). Recently, LLM-based agents show emergent reasoning, planning (liu2024image; achiam2023gpt; anthony_therrien_2024), multimodal understanding (e.g., BLIP-2 (1996cs........5103K)), dynamic task decomposition (e.g., Voyager (sivakorn2016robot)), and tool use (e.g., Toolformer (10.1023/A:1015008417172)), enabling general-purpose intelligence (russell2021modern; song2024anti). Their zero-shot generalization (sumers2024cognitivearchitectureslanguageagents) and social collaboration abilities (1571980075547310080) mark a paradigm shift in agent research.
Advances and Challenges in CAPTCHA Security
Early text CAPTCHAs relied on heavy distortion and noise, but CNN-based segmentation attacks soon defeated many schemes, including reCAPTCHA (chellapilla2005building; shet2014you; bursztein2011text). Image-based CAPTCHAs, such as ASIRRA, emphasized visual cognition yet were ultimately bypassed by SVM classifiers trained on public datasets (ding2025illusioncaptcha; gao2021research). Modern defenses now employ deep object detection, diffusion models, and style transfer for increased complexity (mann2020language; van2023anti; 109c6f85f2734b649d5f7b7a1946b2c1; ci2024ringid; 10.5555/2974989; NEURIPS2023_6dcf277e). Interactive CAPTCHAs (e.g., reCAPTCHA v2) add user gestures like clicking or dragging to resist automation (10.5555/1625995.1626091). Nonetheless, challenges remain in adaptive difficulty, device consistency, adversarial robustness, and reproducible benchmarks (jiang2023diff; alsuhibany2016benchmark).
Model Pass@k
3×3
grid sel.
3×3
jig-swap
Arith.
char.
Arith.
sel.
Hollow
pattern
Distort.
word
Classic
char.
Sequential
letter
Bright.
dist.
Sliding
block
ChatGPT-4o Pass@2 0.780 0.010 0.360 0.015 0.980 0.920 0.045 0.840 0.520 0.220
Pass@3 0.780 0.010 0.360 0.020 0.985 0.920 0.050 0.840 0.520 0.225
Pass@4 0.780 0.010 0.360 0.020 0.985 0.920 0.050 0.840 0.520 0.225
Pass@5 0.780 0.015 0.360 0.020 0.990 0.920 0.050 0.840 0.520 0.225
Seed1.5-VL Pass@2 0.800 0.005 0.320 0.020 0.865 0.960 0.085 0.820 0.465 0.185
Pass@3 0.800 0.005 0.320 0.020 0.865 0.960 0.085 0.825 0.465 0.190
Pass@4 0.805 0.005 0.320 0.020 0.870 0.960 0.085 0.825 0.465 0.195
Pass@5 0.805 0.005 0.320 0.020 0.875 0.960 0.085 0.825 0.465 0.195
Gemini2.5-Pro Pass@2 0.820 0.005 0.360 0.015 0.940 0.920 0.140 0.905 0.480 0.230
Pass@3 0.825 0.005 0.360 0.015 0.945 0.920 0.140 0.905 0.480 0.235
Pass@4 0.825 0.010 0.360 0.020 0.945 0.920 0.145 0.910 0.485 0.235
Pass@5 0.825 0.010 0.360 0.020 0.945 0.920 0.150 0.910 0.485 0.235
Original Qwen-2.5vl-7B-Instr. Pass@2 0.660 0.010 0.365 0.010 0.940 0.020 0.025 0.340 0.220 0.100
Pass@3 0.660 0.010 0.365 0.010 0.940 0.020 0.025 0.340 0.220 0.100
Pass@4 0.665 0.015 0.365 0.010 0.950 0.025 0.030 0.345 0.225 0.100
Pass@5 0.665 0.015 0.365 0.010 0.950 0.025 0.030 0.345 0.225 0.100
Human Pass@2 0.960 0.980 0.990 0.955 0.995 0.885 0.930 0.960 0.965 0.980
Pass@3 0.985 0.985 0.990 0.955 0.995 0.965 0.950 0.980 0.985 0.985
Pass@4 0.995 0.985 0.995 0.965 1.000 0.985 0.980 0.985 0.985 0.990
Pass@5 1.000 0.990 1.000 0.970 1.000 0.990 0.995 0.995 0.990 0.995
Finetuned Qwen-2.5vl-7B-Instr. Pass@2 0.965 0.015 0.700 0.025 0.995 0.985 0.325 0.925 0.665 0.365
Pass@3 0.970 0.015 0.705 0.025 0.995 0.985 0.325 0.925 0.665 0.365
Pass@4 0.975 0.015 0.705 0.030 0.995 0.985 0.325 0.925 0.670 0.365
Pass@5 0.980 0.020 0.710 0.035 1.000 0.995 0.330 0.930 0.675 0.365
Model Pass@k
Align.
sliders
Rotate
block
Geom.
shape
Rotat.
letter
Color
discr.
Vowel
sel.
Full-img
grid sel.
Text-based
arith.
Common
sense
Invert.
letter
ChatGPT-4o Pass@2 0.420 0.140 0.440 0.100 0.960 0.540 0.240 0.980 0.925 0.425
Pass@3 0.420 0.140 0.440 0.100 0.960 0.540 0.240 0.985 0.925 0.425
Pass@4 0.420 0.140 0.440 0.100 0.960 0.540 0.245 0.985 0.925 0.430
Pass@5 0.420 0.140 0.440 0.100 0.960 0.540 0.245 0.985 0.930 0.430
Seed1.5-VL Pass@2 0.380 0.180 0.840 0.140 0.930 0.880 0.230 0.905 0.840 0.360
Pass@3 0.380 0.180 0.840 0.140 0.930 0.880 0.235 0.910 0.845 0.365
Pass@4 0.380 0.185 0.840 0.140 0.930 0.880 0.235 0.920 0.845 0.365
Pass@5 0.380 0.185 0.845 0.140 0.935 0.880 0.235 0.920 0.850 0.365
Gemini2.5-Pro Pass@2 0.420 0.160 0.845 0.120 0.925 0.900 0.280 0.880 0.860 0.385
Pass@3 0.420 0.160 0.845 0.120 0.925 0.900 0.280 0.885 0.865 0.385
Pass@4 0.420 0.165 0.845 0.120 0.925 0.905 0.285 0.890 0.870 0.390
Pass@5 0.420 0.165 0.850 0.125 0.930 0.905 0.290 0.895 0.875 0.400
Original Qwen-2.5vl-7B-Instr. Pass@2 0.300 0.060 0.885 0.100 0.700 0.760 0.140 0.960 0.920 0.365
Pass@3 0.305 0.060 0.885 0.100 0.700 0.760 0.140 0.960 0.920 0.365
Pass@4 0.305 0.060 0.885 0.100 0.700 0.760 0.140 0.965 0.925 0.365
Pass@5 0.305 0.065 0.885 0.100 0.700 0.760 0.145 0.965 0.925 0.370
Human Pass@2 0.975 0.965 0.990 0.875 0.970 0.960 0.940 0.970 0.875 0.955
Pass@3 0.995 0.990 0.990 0.950 0.975 0.980 0.975 0.985 0.905 0.985
Pass@4 1.000 0.995 0.995 0.985 0.990 0.995 0.990 0.990 0.955 1.000
Pass@5 1.000 0.995 1.000 0.990 0.995 1.000 0.995 1.000 0.985 1.000
Finetuned Qwen-2.5vl-7B-Instr. Pass@2 0.565 0.285 0.960 0.335 0.990 0.975 0.360 0.985 0.975 0.520
Pass@3 0.565 0.285 0.960 0.335 0.990 0.975 0.360 0.985 0.975 0.520
Pass@4 0.565 0.285 0.960 0.335 0.995 0.975 0.370 0.990 0.980 0.525
Pass@5 0.570 0.285 0.960 0.340 0.995 0.980 0.370 0.995 0.990 0.530
Table 1: Performance of vision-language models on CAPTCHA-Bench. This table presents Pass@k accuracy on 20 CAPTCHA task types, comparing models and human performance.
@@
Figure 2: Schematic overview of the MCA-Bench data-acquisition and annotation workflow. From left to right, the four grey panels are Static Visual Recognition, Interactive Manipulation, Point-and-Click Localization, and Textual Logic Q&A; the red labels mark each category’s data-collection pipeline. Each pipeline has four stages: (i) define the raw input format; (ii) apply task-specific geometric transforms, coordinate projections, or prompt/template generation; (iii) separate fine-grained annotation types; and (iv) save the annotations to text files.
Figure 3: Schematic of Data Flow Across Four Framework Stages. The schematic diagram illustrates the data flow and key module configuration across the four stages of the end-to-end framework: unified interface access, gent fine-tuning loading, collaborative inference execution, structured result feedback.
Figure 4: Performance Comparison of Multimodal Language Models on MCA-Bench CAPTCHA Tasks. The figure compares the success rates of models including Qwen2.5-VL-7B-Instruct, ChatGPT-4o, Seed1.5-VL, Gemini2.5-Pro, and fine-tuned Qwen2.5-VL-7B-Instruct across MCA-Bench CAPTCHA tasks, covering basic visual recognition, character-based recognition, and advanced multi-step reasoning challenges. Results show fine-tuning consistently improves performance, yet even top-performing models lag behind human-level robustness in complex reasoning tasks.
Dataset Construction Process
@@
This section introduces the MCA Bench pipeline—from raw sample collection to final release—detailing processing and annotation strategies across four task clusters and twenty subtasks. See Figure 2 for the complete workflow.
Data Collection Sources and Compliance Strategies
@@
In developing the MCA-Bench benchmark suite, we implemented a comprehensive data collection and compliance management system to uphold scientific integrity and ensure legal and ethical standards in evaluating CAPTCHA security. Our approach combines independently created data and reused public datasets. For the former, we designed a diverse CAPTCHA dataset, including distorted text, obfuscated graphics, interactive tasks, and puzzles, based on a thorough analysis of technology trends and security needs. These designs are original, validated, and proprietary to our team. For reused datasets, we carefully selected authorized academic, open-source, and industry resources, adhering to licensing agreements and privacy regulations. We ensured compliance through preprocessing steps like anonymization and mitigating privacy risks. Additionally, we established a multi-level review system with legal and ethics oversight, encrypted storage, and tiered access control to guarantee ongoing compliance and data security for MCA-Bench.
Data Collection and Processing
Text-based Task Data Collection
@@
For text-based CAPTCHA tasks, we designed a semi-automated, LLM-driven pipeline to efficiently generate and filter math and commonsense questions, minimizing manual effort while preserving quality and diversity. Using adaptive prompts, the Qwen LLM produced structured, semantically relevant QA pairs, guided by knowledge constraints, task-aware sampling, and type-controlled difficulty. Outputs were refined via manual filtering with a custom evaluation protocol assessing grammar, reasoning, and ambiguity to ensure clarity and robustness for real-world deployment.
Click-based Coordinate Task Data Collection
@@
We propose CPTL (Click-based Positioning and Target Localization), a multimodal benchmark designed to evaluate models’ spatial localization and image-language alignment across varying complexities. It consists of two tasks: Free-form Pointing and Grid-based Discrete Pointing. In Free-form Pointing, we combine procedural background perturbations with public datasets, such as Flickr scenic photos, to create images rich in semantic content. The Grid-based Discrete Pointing task uses a 3×3 grid to evaluate decision-making within specific regions, dividing the image into 9 segments (0–8). The image content includes a 64-class animal dataset, Flickr scenic backgrounds, Google ReCaptcha V2 challenges, and manually designed targets.
@@
CPTL enables adjustable evaluation of spatial complexity and instruction modalities, effectively isolating spatial parsing from language comprehension. It simulates real-world CAPTCHA conditions, testing model robustness against noisy backgrounds, multi-target interference, and weak prompts. Experiments demonstrate that models fine-tuned with CPTL achieve strong localization and semantic understanding, establishing a reliable benchmark for future CAPTCHA security research.
Static Visual Recognition Task Data Collection
We propose SVT (Static Visual Textual Understanding), a benchmark to evaluate multimodal models’ ability to recognize and reconstruct text from distorted images. SVT leverages procedural image generation and linguistic constraints, applying character-level geometric distortions, stochastic noise, and occlusions to create challenging yet interpretable samples. It assesses models’ fine-grained visual attention and robustness to structural noise, spanning from perception to symbolic reasoning. Experiments demonstrate SVT’s effectiveness in revealing limitations of pretrained multimodal models in character-level understanding.
Interactive Behavior Task Data Collection
@@
We propose a multimodal interaction modeling framework for interactive CAPTCHA tasks, leveraging real user data to enhance vision-language models’ dynamic understanding. Focusing on continuous motion interactions—such as sliding alignment, rotation calibration, and trajectory restoration—the framework requires models to interpret spatial structures, motion directions, and behavioral patterns. We collected diverse user trajectories, annotated with timestamps and velocity, enabling realistic spatiotemporal modeling.
@@
Tasks are structured around target-state restoration, divided into standardized subtasks with clear start, goal, and intermediate states. Interaction trajectories are serialized as behavioral vectors combined with visual data for joint learning, significantly enhancing action alignment. To evaluate models comprehensively, we propose new metrics: Center Deviation Error, Angular Restoration Accuracy, Slide Path Alignment Rate, and Motion Variability Index.
Data Annotation Strategy